OpenStudy (anonymous):

In what quadrant will θ terminate, if sin θ is positive and sec θ is negative? Quadrant I Quadrant II Quadrant III Quadrant IV

OpenStudy (campbell_st):

sec = 1/cos for both sin and cos to be positive you are in the 1st quadrant

OpenStudy (anonymous):

2nd.

OpenStudy (anonymous):

what? 1st or 2nd?

OpenStudy (ash2326):

\(\sin \theta\) is positive in two quadrants I and II In I= \(\cos\theta \) is positive In II= \(\cos\theta \) is negative so it's second quadrant

OpenStudy (anonymous):

2nd. You wnt cos negative sin positive. Thats the scond.

OpenStudy (campbell_st):

 the picture shows which trig ratios are positive in which quadrants